# Robust Spatial Confounding Adjustment via Basis Voting

### v1

@@ 3 Plurality Rule Mode Estimator
In analogy to instrumental variables, a basis function $h_{j}$ is appropriate for constructing a candidate estimator $\hat{\beta}_{j}$ if it captures variation in the exposure but not in the unmeasured confounder. Such a direction isolates unconfounded exposure variation, enabling consistent estimation of the exposure’s effect on the outcome. In the basis representation of the exposure $X$ and the unmeasured confounder $U$ in Section 2, projections of a basis function $h_{j}$ on them are given by the coefficients $\alpha_{x}^{(j)}$ and $\alpha_{u}^{(j)}$. Thus a basis function $h_{j}$ is an ingredient of a valid candidate if $j\in O_{X}$, i.e., $\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=0$. On the other hand $h_{j}$ (and, $\hat{\beta}_{j}$) will be invalid if $\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=\nu$ for some invalidity level $\nu\neq 0$. Causal methods handle multiple IVs of unknown validity based on assumptions about the number of valid instruments. The majority rule assumes that more than $50\%$ of the candidate instruments are valid, in which case $\beta$ can be identified by simply using the median of the candidate instrument-based estimators. In the context of spatial confounding, saying that the confounder is not supported by over half the basis set is a strong assumption.
Plurality rule clusters the basis functions $h_{j}$ in terms of the ratio $\nu$, and posits that the biggest cluster is at 0. Of course, if majority rule holds then the cluster at 0 includes more than $50\%$ of the basis, and plurality rule automatically holds. Hence, plurality is a weaker assumption than majority. In fact, plurality does not enforce any minimal cardinality of $O_{X}$, which can be as low as $2$ (if the remaining ratios are all unique) as instead of the median we will use a mode based estimator. We define the plurality rule formally below. Note that we will denote in-probability convergence using $\xrightarrow[n\to\infty]{P}$ and weak convergence using $\xrightarrow[n\to\infty]{d}$ throughout.
@@
We define the set $S_{x,d}=\left\{1\leq k\leq d:\alpha^{(k)}_{x}\neq 0\right\}$, that is, the set denoting the indices of the basis functions which have non-zero projection on $X$. This set is identifiable because we observe $X$ and we can find the projection $\alpha_{x}^{(k)}$ of $X$ on each $h_{k}$ filtering it based on a suitable small threshold to determine which $k$ lies in $S_{x,d}$. Next, we define the set of confounder-exposure projection ratios for the basis functions with indices in the set $S_{x,d}$ as:
$R_{d}=\left\{r_{k,d}\coloneqq\dfrac{\alpha_{u}^{(k)}}{\alpha_{x}^{(k)}}:k\in S_{x,d}\text{ i.e }\alpha_{x}^{(k)}\neq 0\right\}$ (2)
@@
Let $\mathcal{C}_{n}=\left\{\hat{\beta}_{k,n}:k\in S_{x,d}\right\}$ be a class of candidate estimators, where $\hat{\beta}_{k,n}$ corresponds to the basis function $h_{k}$ such that for each $k\in S_{x,d}$, $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{k,d}$. Specific candidate estimators will be presented in Section 4, but the main idea of the plurality-rule based estimator that we discuss in this section is agnostic to this choice. For any $\nu\in\mathbb{R}$, define
$\mathcal{I}_{\nu}=\left\{r_{k,d}\in R_{d}:r_{k,d}=\nu\right\}$ (3)
@@
Then, $k\in O_{X}\iff r_{k,d}\in\mathcal{I}_{0}$ and this one-to-one correspondence gives us the rationale for the next assumption, which will help us estimate $\beta$ without the knowledge of $O_{X}$:
Assumption 2 (Plurality Assumption).
$\left|\mathcal{I}_{0}\right|>\underset{\nu\neq 0}{\max}\left|\mathcal{I}_{\nu}\right|$.
@@
Under this assumption, the cluster of ratios $r_{k,d}$ at 0 is larger than clusters at any other real number $\nu$, implying that the mode of the set $R_{d}$ is 0. To illustrate this is a reasonable assumption in real world applications, we chose the TerraClimate dataset (Abatzoglou et al. (2018)), which provides monthly climate variables globally at $\sim 4$ km spatial resolution. Specifically, we extracted two meteorological variables, namely monthly minimum temperature $(^{\circ}C)$ and total precipitation (mm), for the year 2024, averaged them over the 12 months, and restricted the domain to the contiguous United States. These variables have an empirical correlation of 0.42, and in an analysis either of them can play the role of an exposure or an unmeasured confounder. Here, we have considered a set of basis functions that are the cartesian product of two one-dimensional Fourier basis functions and thus lack the natural roughness ordering among them. These annual mean fields were then projected onto this basis function set and we looked at both the basis coefficients as well as the mode of $R_{d}$ for varying choices of $d$. Figure 3 presents the illustration.
@@
Figure 3: (a) Annual mean precipitation (Ppt) in $m$ and (b) annual mean minimum temperature (Tmin) in $^{\circ}C$ for 2024 over the contiguous U.S. Absolute Fourier basis coefficients for (c.1) Ppt and (c.2) Tmin using 10,000 basis functions. Absolute density mode of coefficient ratios: (d) Ppt/Tmin and (e) Tmin/Ppt across varying numbers of basis functions.
@@
We see from comparing Panels (a) and (b) that the map of temperature looks much smoother than that of precipitation. This is also confirmed by the basis coefficient plots in panels (c.1) and (c.2) showing that the support of precipitation has a much longer tail. Existing spatial deconfounding methods would thus do well when the precipitation is the exposure and the temperature is the unmeasured confounder but not the other way around. To assess whether the plurality rule holds in either direction, we plot the modes of $R_{d}$ as a function of $d$ both ways – with precipitation as confounder and temperature as exposure in Panel (d), and vice versa in Panel (e). We see that in both plots, the modes decay to 0 as $d$ increases, suggesting that the plurality rule is satisfied in either direction, and estimators based on the plurality rule can account for a wider range of scenarios of spatial confounding.
@@
We formalize the estimation below. As $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{k,d}$, if the mode of $R_{d}$ is 0, intuitively the mode of $\mathcal{C}_{n}$ will be $\beta+\text{mode}(R_{d})=\beta$. Thus the mode of the candidate estimators can be a good estimator for $\beta$. In practice, as $\mathcal{C}_{n}$ is a finite set, it is likely that each value of $\hat{\beta}_{k,n}$ will be unique. So we will use a kernel density smoothing to estimate the mode.
@@
The following Lemma shows that the mode of a KDE based on a finite set of values with unique mode also has the same mode with a suitably chosen bandwidth.
Lemma 1.
If $K$ is Lipschitz continuous and the finite set $\bm{a}_{m}=\{a_{1},\cdots,a_{m}\}$ has a unique mode $t$, then for some $h_{0}$ and all $0<h\leq h_{0}$, where $h_{0}$ depends on $K$ and $\underset{a_{i}\neq t}{\min}|t-a_{i}|$, $\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\bm{a}_{m},K,h,x)=t$.
Hence, we work with the KDE estimate for a small enough bandwidth which depends on the functional form of the kernel and the separation of 0 from any other mode of the set $R_{d}$. As the limit points of $\mathcal{C}_{n}$ are $\{\beta+r_{k,d}:r_{k,d}\in R_{d}\}$ we then have, $\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\beta+R_{k,d},K,h,x)=\beta+\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(R_{k,d},K,h,x)=\beta$ where $\beta+R_{d}=\{\beta+r_{k,d}:r_{k,d}\in R_{d}\}$. This motivates us to propose the following ensemble estimator by combining all the candidate estimators from the class $\mathcal{C}_{n}$:
$\hat{\beta}_{n}=mode(\mathcal{C}_{n})\coloneqq\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\mathcal{C}_{n},K,h,x).$ (4)
We call our method ‘basis voting’ as the process to select the final estimator $\hat{\beta}_{n}$ mimics a voting process. Each basis $h_{k}$ represents a voter who votes for a candidate $\hat{\beta}_{k,n}$, and the most popular candidate (formally, the KDE mode) among the set of candidates $\mathcal{C}_{n}=\{\hat{\beta}_{k,n}\}$ is the winner (final estimator). The following theorem gives us theoretical guarantees for the mode estimator:
@@
Consider the class of estimators $\mathcal{C}_{n}=\left\{\hat{\beta}_{k,n}:k\in S_{x,d}\right\}$ such that for each $k\in S_{x,d}$, $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{k,d}$ where $r_{k,d}\in R_{d}$ defined in Equation 2, then under Assumption 2, a Lipschitz continuous Kernel $K$, and bandwidth $h\leq h_{0}$ as defined in Lemma 1,
$\hat{\beta}_{n}=\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\mathcal{C}_{n},K,h,x)\xrightarrow[n\to\infty]{P}\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\{\beta+r_{k,d}:r_{k,d}\in R_{d}\},K,h,x)=\beta.$
Hence, we have a consistent estimator of $\beta$ under the plurality rule without any assumptions on the exposure being rougher than the confounder. As we see, the consistency of the ensemble model estimator relies on consistency of candidate estimators $\hat{\beta}_{k,n}$ to the respective limits $\beta+r_{k,d}$. In order to get good, efficient estimates of $\beta$, we need to have our candidate estimators be well-behaved and have efficient rates of convergence towards their limits. The following section presents choices of such candidate estimators.
@@ 4 Candidate Estimators
The building blocks of the mode estimator $\hat{\beta}_{n}$ are the candidate estimators $\hat{\beta}_{j,n}$ and thus we need to have efficient choices for them. In this section, we present two sets of candidates. We start with the classical projection-based estimator commonly used in the causal inference literature, and then we will introduce a new class of candidate estimators, comparing its performance with the former one. Our goal in this section is to demonstrate how our new candidate estimates can improve on both bias and efficiency in estimating the causal effect $\beta$. We define some notations that we will be using to define our candidate estimators. For a given set of sampled spatial coordinates $(S_{1},\dots,S_{n})$, let $\bm{h}_{n,j}=(h_{j}(S_{1}),\dots,h_{j}(S_{n}))^{\top}$ denote the vector of evaluated values for the basis $h_{j}$. Also let us denote $\bm{H}_{1:k}=\begin{bmatrix}\bm{h}_{n,1}&\cdots&\bm{h}_{n,k}\end{bmatrix}$ the evaluation matrix for first $k$ basis functions and corresponding projection on its column space as $\bm{P}_{k}=\bm{H}_{1:k}\left(\bm{H}_{1:k}^{\top}\bm{H}_{1:k}\right)^{-1}\bm{H}_{1:k}^{\top}$. Define the projected exposures and outcomes on the first $d$ basis as $\bm{X}_{n,d}=\bm{P}_{d}\bm{X}_{n}$ and $\bm{Y}_{n,d}=\bm{P}_{d}\bm{Y}_{n}$.
@@ 4.1 Projection-Based Estimator
We begin with a standard projection-based candidate estimator that aligns closely with classical approaches. In the context of time-series deconfounding, this candidate has been used by (Schur & Peters 2024) under a special case of the plurality-type assumption. Consider any single basis function $h_{j}(\cdot)$ with index $j\in S_{x,d}$, indicating that $h_{j}$ contributes to the spatial variation in the exposure $X$. We use this basis function to define a classic Two-Stage Least Squares (2SLS) estimator. The projection matrix onto $\bm{h}_{n,j}$ is given by $\bm{P}_{n,j}\coloneqq\bm{h}_{n,j}\left({\bm{h}_{n,j}^{\top}\bm{h}_{n,j}}\right)^{-1}\bm{h}_{n,j}^{\top}.$ Then, for any $j\in S_{x,d}$, the projection candidate is computed as the OLS between the projections of the outcome and the exposure, i.e.,
$\hat{\beta}_{j,n}^{PROJ}\coloneqq\left[\left(\bm{P}_{n,j}\bm{X}_{n}\right)^{\top}\left(\bm{P}_{n,j}\bm{X}_{n}\right)\right]^{-1}\left(\bm{P}_{n,j}\bm{X}_{n}\right)^{\top}\bm{P}_{n,j}\bm{Y}_{n}$ (5)
@@
Under the DGP as in Equation 1 and the estimator as in Equation 5, assuming a sampling density $f_{S}>0$ on the compact domain $\mathcal{D}$ for the i.i.d. sampled coordinates with $\epsilon(S_{i})|S_{i}$’s being i.i.d with mean 0 and variance $\sigma^{2}_{\epsilon}$, defining $r_{j,d}=\alpha_{u}^{(j)}/\alpha_{x}^{(j)}$ such that $j\in S_{x,d}$, we have:
@@
    $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta+r_{j,d}$
@@
    $\sqrt{n}\left(\hat{\beta}_{j,n}^{PROJ}-(\beta+r_{j,d})\right)\xrightarrow[n\to\infty]{d}N\left(0,\sigma^{2}_{j,d}\right)$
    where $\sigma^{2}_{j,d}\coloneqq\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}+\alpha_{u}^{(j)}\left(\frac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{4}}\sigma^{2}_{X,j}-\frac{2}{{\alpha_{x}^{(j)}}^{3}}\sigma_{U,X,j}\right)$ for $\sigma^{2}_{U,j}\coloneqq\|h_{j}U\|^{2}_{f_{S}}-{\alpha_{u}^{(j)}}^{2}$, $\sigma^{2}_{X,j}\coloneqq\|h_{j}X\|^{2}_{f_{S}}-{\alpha_{x}^{(j)}}^{2}$ and $\sigma_{U,X,j}\coloneqq\langle h_{j}X,h_{j}U\rangle_{f_{S}}-{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}$.
The theorem proves that the projection-based candidate estimators $\hat{\beta}_{j,n}\xrightarrow[n\to\infty]{P}\beta+r_{j,d}$ thereby satisfying the requirement to be used for the mode estimator (4). It also provides the asymptotic variance of the candidates, which, as we will show, is important in determining efficiency. An immediate corollary from the theorem is the consistency and asymptotic normality of the valid projection candidates.
@@
Under the setup mentioned in Theorem 2, for any $j\in O_{X}$ defined in Assumption 1, we have $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta$, and after proper scaling, we have the asymptotic distribution $\sqrt{n}\left(\hat{\beta}_{j,n}^{PROJ}-\beta\right)\xrightarrow[n\to\infty]{d}N\left(0,\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}\right)$ where $\sigma^{2}_{U,j}=\|h_{j}U\|^{2}_{f_{S}}$.

### v2

@@ 3 Plurality Rule Mode Estimator
We say that a basis function $h_{j}$ is valid if it gives rise to a candidate estimator $\hat{\beta}_{j}$ that is consistent for $\beta$. In order to be valid, the basis function must capture variation in the exposure but not in the unmeasured confounder, i.e., it must isolate unconfounded exposure variation. This is the case when $j\in O_{X}$, i.e., $\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=0$. On the other hand $h_{j}$ (and, $\hat{\beta}_{j}$) will be invalid if $\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=\nu$ for some value $\nu\neq 0$.
The plurality rule assumes that invalid estimators tend to have distinct biases and probability limits, and therefore in the set of probability limits of all candidate estimators, the largest cluster, i.e., the plurality, corresponds to the set of valid estimators. In our case, we will show that this assumption clusters the basis functions $h_{j}$ in terms of the ratio $\nu$, and posits that the biggest cluster is at 0. The plurality rule does not enforce any minimal cardinality of $O_{X}$ in terms of the sample size. In fact, $|O_{X}|$ can be as low as $2$ (provided no two of the remaining ratios are equal). We define the plurality rule formally below.
@@
For any $d^{\prime}\in\mathbb{N}$, define the set $S_{X,d^{\prime}}=\left\{1\leq k\leq d^{\prime}:\alpha^{(k)}_{x}\neq 0\right\}$, that is, the set denoting the indices of the basis functions that have non-zero projection on $X$. Next, define the set of confounder-exposure projection ratios for the basis functions with indices in the set $S_{X,d^{\prime}}$ as:
$R_{d^{\prime}}=\left\{r_{k}\coloneqq\dfrac{\alpha_{u}^{(k)}}{\alpha_{x}^{(k)}}:k\in S_{X,d^{\prime}}\text{ i.e }\alpha_{x}^{(k)}\neq 0\right\}$ (2)
@@
Let $\mathcal{C}_{n,d^{\prime}}=\left\{\hat{\beta}_{k,n}:k\in S_{X,d^{\prime}}\right\}$ be a class of candidate estimators, where $\hat{\beta}_{k,n}$ corresponds to the basis function $h_{k}$ such that for each $k\in S_{X,d^{\prime}}$, $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{k}$. We denote in-probability convergence using $\xrightarrow[n\to\infty]{P}$ and weak convergence using $\xrightarrow[n\to\infty]{d}$ throughout. Specific candidate estimators will be presented in Section 4, but the main idea of the plurality-rule based estimator is agnostic to this choice. For any $\nu\in\mathbb{R}$, define
$\mathcal{I}_{\nu,d^{\prime}}=\left\{r_{k}\in R_{d^{\prime}}:r_{k}=\nu\right\}$ (3)
@@
Then, $k\in O_{X,d^{\prime}}\iff r_{k}\in\mathcal{I}_{0,d^{\prime}}$ and this one-to-one correspondence gives us the rationale for the next assumption, which will help us estimate $\beta$ without the knowledge of $O_{X}$:
Assumption 2 (Plurality Rule).
There exists some $d_{0}\in\mathbb{N}$, such that for all $d^{\prime}\geq d_{0}$, $\left|\mathcal{I}_{0,d^{\prime}}\right|>\left|\mathcal{I}_{\nu,d^{\prime}}\right|$ for any $\nu\neq 0$.
@@
The plurality rule posits that when considering a sufficiently large number of basis functions $h_{k}$, the cluster of ratios $r_{k}$ at 0 is larger than clusters of these ratios at any other real number $\nu$. This implies that when using $d\geq d_{0}$ of basis functions for the analysis, the mode of the set $R_{d}$ is 0. To illustrate that this is a reasonable assumption in real world applications, we extracted two meteorological variables–monthly minimum temperature $(^{\circ}C)$ and total precipitation (mm), for the year 2024– from the TerraClimate dataset (abatzoglou2018terraclimate), which provides monthly climate variables globally at $\sim 4$ km spatial resolution. We averaged them over the 12 months and restricted the domain to the contiguous United States. These variables have an empirical correlation of 0.42, and in an analysis either of them can play the role of an exposure or an unmeasured confounder. Here, we have considered a set of basis functions that are the Cartesian product of two one-dimensional Fourier basis functions and thus lack a natural frequency ordering. We projected both variables onto this basis function set and looked at both the basis coefficients as well as the mode of $R_{d}$ for varying choices of $d$. Figure 3 presents the illustration.
@@
Figure 3: (a) Annual mean precipitation (Ppt) in $m$ and (b) annual mean minimum temperature (Tmin) in $^{\circ}C$ for 2024 over the contiguous U.S. Absolute Fourier basis coefficients for (c.1) Ppt and (c.2) Tmin using 10,000 basis functions. Mode of the density of absolute values of the coefficient ratios: (d) Ppt/Tmin and (e) Tmin/Ppt across varying numbers of basis functions.
@@
Comparing Panels (a) and (b), that the map of temperature looks much smoother than that of precipitation. This is confirmed by the basis coefficient plots in panels (c.1) and (c.2), which show that the support of precipitation has a longer tail. Existing spatial deconfounding methods would thus do well when precipitation is the exposure and temperature is the unmeasured confounder, but not the other way around. To assess whether the plurality rule holds in either direction, we plot the modes of $R_{d}$ as a function of $d$ both ways – with precipitation as confounder and temperature as exposure in Panel (d), and vice versa in Panel (e). We see that in both plots, the modes decay to 0 and stabilize as $d$ increases, suggesting that the plurality rule is satisfied in either direction, and estimators based on the plurality rule can account for a wider range of scenarios of spatial confounding.
@@
We formalize the estimation of the mode below. As $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{d}$, if the mode of $R_{d}$ is 0, intuitively the mode of $\mathcal{C}_{n,d}$ will be $\beta+\text{mode}(R_{d})=\beta$. Thus the mode of the candidate estimators can be a good estimator for $\beta$. In practice, it is likely that each value of the estimates $\hat{\beta}_{k,n}$ will be unique, and we use a kernel density smoothing to estimate the mode.
@@
The mode of a KDE based on a finite set of values with unique mode also has the same mode, for small enough bandwidth (See Appendix S3.4). Here, the bandwidth will depend on the functional form of the kernel $K$ and the separation from 0 of any other cluster of the set $R_{d}$. As the limit points of $\mathcal{C}_{n,d}$ are $\beta+R_{d}:=\{\beta+r_{k}:r_{k}\in R_{d}\}$, using shift-invariance of mode [^1] If $X$ is a random variable with mode at $\mu$, then for any $c\in\mathbb{R}$, $X+c$ has mode at $\mu+c$., we then have $\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\beta+R_{d},K,h,x)=\beta+\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(R_{d},K,h,x)=\beta$. This motivates us to propose the following ensemble estimator by combining all the candidate estimators from the class $\mathcal{C}_{n,d}$:
$\hat{\beta}_{n}=\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\mathcal{C}_{n,d},K,h,x).$ (4)
We call our method ‘basis voting’ as the process to select the final estimator $\hat{\beta}_{n}$ mimics a voting process. Each basis $h_{k}$, $k=1,\ldots,d$, represents a voter who votes for a candidate $\hat{\beta}_{k,n}$, and the most popular candidate (formally, the KDE mode) among the set of candidates $\mathcal{C}_{n,d}=\{\hat{\beta}_{k,n}\}$ is the winner, i.e., the final estimator. The following theorem gives us theoretical guarantees for the mode estimator. All proofs are in the supplement.
@@
Consider the class of estimators $\mathcal{C}_{n,d}=\left\{\hat{\beta}_{k,n}:k\in S_{X,d}\right\}$ such that for each $k\in S_{X,d}$, $\hat{\beta}_{k,n}\xrightarrow[n\to\infty]{P}\beta+r_{k}$ where $r_{k}\in R_{d}$ defined in Equation 2. Then under Assumption 2, for any $d\geq d_{0}$, a Lipschitz continuous Kernel $K$, and bandwidth $h\leq h_{0}$ where $h_{0}$ is some constant depending on $K$ and $\underset{r_{k}\in R_{d}}{\min}\left|r_{k}\right|$, with $\hat{\beta}_{n}$ as defined in (4), we have $\hat{\beta}_{n}\xrightarrow[n\to\infty]{P}\beta$.
From Theorem 1, we have a consistent estimator of $\beta$ under the plurality rule without any assumption about the relative smoothness of the exposure and confounder. The result holds as long as we are using $d\geq d_{0}$ number of basis functions. In practice, as $d_{0}$ is unknown, we recommend repeating the analysis for a range of choice of $d$ and checking that the estimates stabilize after sufficiently large $d$.
The consistency of the ensemble mode-based basis voting estimator relies on the convergence in probability of the candidate estimators $\hat{\beta}_{k,n}$ to their respective limits $\beta+r_{k}$. So, the performance and efficiency of our estimator of $\beta$ requires well-behaved and efficient candidate estimators. The following section presents two such candidate estimators.
@@ 4 Candidate Estimators
The building blocks of the mode estimator $\hat{\beta}_{n}$ are the candidate estimators $\hat{\beta}_{j,n}$ and in this section, we present two sets of candidates. We start with a projection-based estimator similar to two-stage least squares and then we introduce a new drop-one candidate estimator, comparing its performance with the projection-based estimator. Our goal in this section is to demonstrate how our new drop-one candidate estimator can improve on both bias and efficiency in estimating $\beta$.
@@ 4.1 Projection-Based Estimator
We begin with a projection-based candidate, similar to two-stage least squares. In the context of time-series deconfounding, this candidate has been used by (schur2025decor) under a special case of the plurality-type assumption. Consider any single basis function $h_{j}(\cdot)$ with index $j\in S_{X,d}$, indicating that $h_{j}$ contributes to the spatial variation in the exposure $X$. We use this basis function to define a projection-on-projection estimator: For a given set of sampled spatial coordinates $(S_{1},\dots,S_{n})$, let $\bm{h}_{n,j}=(h_{j}(S_{1}),\dots,h_{j}(S_{n}))^{\top}$ denote the vector of evaluated values for the basis $h_{j}$. The projection matrix onto $\bm{h}_{n,j}$ is given by $\bm{P}_{n,j}\coloneqq\bm{h}_{n,j}\left({\bm{h}_{n,j}^{\top}\bm{h}_{n,j}}\right)^{-1}\bm{h}_{n,j}^{\top}.$ Then, for any $j\in S_{X,d}$, the projection candidate is computed as the OLS between the projections of the outcome and the exposure, i.e.,
$\hat{\beta}_{j,n}^{PROJ}\coloneqq\left[\left(\bm{P}_{n,j}\bm{X}_{n}\right)^{\top}\left(\bm{P}_{n,j}\bm{X}_{n}\right)\right]^{-1}\left(\bm{P}_{n,j}\bm{X}_{n}\right)^{\top}\bm{P}_{n,j}\bm{Y}_{n}.$ (5)
@@
Under the DGP as in Equation 1 and the estimator as in Equation 5, assuming a sampling density $f_{S}>0$ on the compact domain $\mathcal{D}$ for the i.i.d. sampled coordinates with $\epsilon(S_{i})|S_{i}$’s being i.i.d with mean 0 and variance $\sigma^{2}_{\epsilon}$, defining $r_{j}=\alpha_{u}^{(j)}/\alpha_{x}^{(j)}$ for $j$ such that $\alpha_{x}^{(j)}\neq 0$ , we have:
@@
    $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta+r_{j}$
@@
    $\sqrt{n}\left(\hat{\beta}_{j,n}^{PROJ}-(\beta+r_{j})\right)\xrightarrow[n\to\infty]{d}N\left(0,\sigma^{2}_{j}\right)$
    where $\sigma^{2}_{j}\coloneqq\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}+\alpha_{u}^{(j)}\left(\frac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{4}}\sigma^{2}_{X,j}-\frac{2}{{\alpha_{x}^{(j)}}^{3}}\sigma_{U,X,j}\right)$ for $\sigma^{2}_{U,j}\coloneqq\|h_{j}U\|^{2}_{f_{S}}-{\alpha_{u}^{(j)}}^{2}$, $\sigma^{2}_{X,j}\coloneqq\|h_{j}X\|^{2}_{f_{S}}-{\alpha_{x}^{(j)}}^{2}$ and $\sigma_{U,X,j}\coloneqq\langle h_{j}X,h_{j}U\rangle_{f_{S}}-{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}$.
The theorem proves that the projection-based candidate estimators $\hat{\beta}_{j,n}\xrightarrow[n\to\infty]{P}\beta+r_{j}$ thereby satisfying the requirement to be used for the mode estimator (4). It also provides the asymptotic variance of the candidates, which, as we will show, is important in determining efficiency. An immediate corollary from the theorem is the consistency and asymptotic normality of the valid projection candidates.
@@
Under the setup mentioned in Theorem 2, for any $j\in O_{X}$ defined in Assumption 1, we have $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta$, and after proper scaling, we have the asymptotic distribution $\sqrt{n}\left(\hat{\beta}_{j,n}^{PROJ}-\beta\right)\xrightarrow[n\to\infty]{d}N\left(0,\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}\right)$ where $\sigma^{2}_{U,j}$ reduces to $\|h_{j}U\|^{2}_{f_{S}}$.